OpenStudy (anonymous):

anyone 10.5x+61.81=-9.5+61.31

OpenStudy (unklerhaukus):

\[10.5x+61.81=-9.5+61.31\]\[10.5x=-9.5+61.31-61.81\] \[x=\frac{-9.5+61.31-61.81}{10.5}\]